OpenStudy (anonymous):

PLEASE HELP!!!!!!! the annual interest rate on a $6000 investment exceeds the interest earned on a $4000 investment by $204. The $6000 is invested at a 0.4% higher rate of interest than the $4000. What is the interest rate of each investment?

OpenStudy (amistre64):

6000(n%) = 4000(m%) + 204 n% = m% + .4

OpenStudy (amistre64):

6000(m%+.4%) = 4000(m%) + 204 6000(m%) + 6000(.4%) = 4000(m%) + 204 6000(m%) -4000(m%) = 205 - 6000(.4%) m%(6000-4000) = 205 - 6000(.4%) m% = 205 - 6000(.004) --------------- 2000

OpenStudy (amistre64):

m% = .0905 m%(100) = .0905(100) m = 9.05% n = 9.05% + .4% = 9.45% dbl chk

OpenStudy (amistre64):

i see a mistake, just a slight one tho; i mistyped 204 into a 205 ....

OpenStudy (amistre64):

m = 9% n = 9.4% then ... sooo close :)

OpenStudy (amistre64):

that checks out :)